On July 1, 2018, Truman Company acquired a 70 percent interest in Atlanta Company in exchange for consideration of $838,075 in cash and equity securities. The remaining 30 percent of Atlanta’s shares traded closely near an average price that totaled $359,175 both before and after Truman’s acquisition.

In reviewing its acquisition, Truman assigned a $137,000 fair value to a patent recently developed by Atlanta, even though it was not recorded within the financial records of the subsidiary. This patent is anticipated to have a remaining life of five years.

The following financial information is available for these two companies for 2018. In addition, the subsidiary’s income was earned uniformly throughout the year. The subsidiary declared dividends quarterly.

Truman Atlanta
Revenues $ (711,290 ) $ (550,000 )
Operating expenses 413,000 372,000
Income of subsidiary (52,710 ) 0
Net income $ (351,000 ) $ (178,000 )
Retained earnings, 1/1/18 $ (896,000 ) $ (598,000 )
Net income (above) (351,000 ) (178,000 )
Dividends declared 170,000 70,000
Retained earnings, 12/31/18 $ (1,077,000 ) $ (706,000 )
Current assets $ 357,715 $ 531,000
Investment in Atlanta 866,285 0
Land 481,000 234,000
Buildings 789,000 651,000
Total assets $ 2,494,000 $ 1,416,000
Liabilities $ (917,000 ) $ (390,000 )
Common stock (95,000 ) (300,000 )
Additional paid-in capital (405,000 ) (20,000 )
Retained earnings, 12/31/18 (1,077,000 ) (706,000 )
Total liabilities and stockholders' equity $ (2,494,000 ) $ (1,416,000 )
  1. How did Truman allocate Atlanta’s acquisition-date fair value to the various assets acquired and liabilities assumed in the combination?

  2. How did Truman allocate the goodwill from the acquisition across the controlling and noncontrolling interests?

  3. How did Truman derive the Investment in Atlanta account balance at the end of 2018?

  4. Prepare a worksheet to consolidate the financial statements of these two companies as of December 31, 2018. At year-end, there were no intra-entity receivables or payables.
Answer #1

Part A

Consideration transferred by Truman

$

838075

Non-controlling interest fair value

359175

Atlanta’s acquisition-date total fair value

$

1197250

Book value of Atlanta

(972000)

Fair value in excess of book value

$

225250

Excess fair value assigned to specific accounts based on fair value

Remaining life

Annual excess

Amortizations

Patent

137000

5 years

27400

Goodwill

$

88250

indefinite

0

Total

27400

Book value = common stock + additional paid in capital + retained earnings 1/1 + (net income *1/2)

= 300000+20000+598000+((178000-70000)*1/2) = 972000

Goodwill = Fair value in excess of book value-Patent

Part B

Controlling interest

Non-controlling interest

Fair value at acquisition date

$

838075

$

359175

Relative fair values of identifiable net assets 70% and 30% of $1109000 (acquisition date book value + patent = net asset fair value)

776300

332700

Goodwill

$

61775

$

26475

Part C

Initial value at acquisition date

$

838075

Truman’s share of Atlanta’s net income for half year

([($178,000-27400 amortization) X ½ year] X 70%)

52710

Dividends 2018 ($70,000 X ½ year X 70%)

(24500)

Investment account balance 12/31/18

$866285

Part D

TRUMAN COMPANY AND SUBSIDIARY ATLANTA COMPANY

Consolidated Worksheet

For the year ending December 31, 2018

Accounts

Truman

Altanta

Consolidated entries

Non controlling interest

Consolidated totals

Debit

Credit

Revenues

$

(711290)

$

(550000)

(S)

275000

$

$

(986290)

Operating expenses

413000

372000

13700

186000

612700

Net income to subsidiary

(52710)

52710

0

Separate company net income

(351000)

(178000)

Consolidated net income

(373590)

Net income attributable to NCI

(22590)

22590

Net income attributable to Truman

(351000)

Retained earnings 1/1

(896000)

(598000)

598000

(896000)

Net income

(351000)

(178000)

(351000)

Dividends declared

170000

70000

59500

10500

170000

Retained earnings 12/31

(1077000)

(706000)

(1077000)

Current assets

357715

531000

888715

Investment in Atlanta

866285

0

24500

890785

0

Land

481000

234000

715000

buildings

789000

651000

1440000

Patent

137000

13700

123300

goodwill

88250

88250

Total assets

2494000

1416000

3255265

Liabilities

(917000)

(390000)

(1307000)

Common stock

(95000)

(300000)

300000

(95000)

Additional paid in capital

(405000)

(20000)

20000

(405000)

Retained earnings 12/31

(1077000)

(706000)

(1077000)

Noncontrolling interest 7/1

359175

(359175)

Noncontrolling interest 12/31

(371265)

(371265)

Total liabilities and equity

(2494000)

(1416000)

1509160

1509160

(3255265)

550000/2 = 275000

27400/2 = 13700

372000/2 = 186000

70000/2*30% =   10500

(70000/2)+(70000/2*70%) = 59500

(70000/2*70%) = 24500